# Cooperative Sheaf Neural Networks

### v1

@@ Abstract
Sheaf diffusion has recently emerged as a promising design pattern for graph representation learning due to its inherent ability to handle heterophilic data and avoid oversmoothing. Meanwhile, cooperative message passing has also been proposed as a way to enhance the flexibility of information diffusion by allowing nodes to independently choose whether to propagate/gather information from/to neighbors. A natural question ensues: is sheaf diffusion capable of exhibiting this cooperative behavior? Here, we provide a negative answer to this question. In particular, we show that existing sheaf diffusion methods fail to achieve cooperative behavior due to the lack of message directionality. To circumvent this limitation, we introduce the notion of cellular sheaves over directed graphs and characterize their in- and out-degree Laplacians. We leverage our construction to propose Cooperative Sheaf Neural Networks (CSNNs). Theoretically, we characterize the receptive field of CSNN and show it allows nodes to selectively attend (listen) to arbitrarily far nodes while ignoring all others in their path, potentially mitigating oversquashing. Our experiments show that CSNN presents overall better performance compared to prior art on sheaf diffusion as well as cooperative graph neural networks.
@@ 1 Introduction
Graph neural networks (GNNs) have become the standard models for an array of sensitive machine learning over networked data, with far-reaching applications in, e.g., physics simulation (Sanchez-Gonzalez et al., 2020), recommender systems (Ying et al., 2018), and molecular modeling (Duvenaud et al., 2015; Gilmer et al., 2017). Most GNNs follow a message-passing paradigm, in which node representations are updated by aggregating features received from neighboring nodes. This approach assumes that every node participates in communication by both sending and receiving information.
A recent research direction focuses on generalizing GNNs and its message-passing paradigm through Sheaf Neural Networks (SNNs), which incorporate additional structure by modeling interactions using cellular sheaves. A cellular sheaf ${\mathcal{F}}$ over an undirected graph associates (i) vector spaces ${\mathcal{F}}(i)$ and ${\mathcal{F}}(e)$, known as stalks, to each vertex $i$ and each edge $e$, and (ii) a linear map $\mathcal{F}_{i\unlhd e}$, known as a restriction map, to each incident vertex-edge pair $i\unlhd e$. These mathematical constructs induce a continuous-time diffusion process over the graph, which we can discretize to derive neural architectures for relational data. Notably, Neural Sheaf Diffusion (NSD) (Bodnar et al., 2022) model learns these restriction maps and are naturally capable of avoiding oversmoothing while handling graphs with heterotrophic characteristics. In this context, the sheaf Laplacian serves as the main operator to propagate information, extending classical graph convolutional networks (Kipf and Welling, 2017).
In parallel, Cooperative GNNs (Finkelshtein et al., 2024) have been proposed as a way to increase the expressiveness of message-passing by relaxing the assumption that nodes must always send or receive information. In this framework, nodes can choose to ignore the features of their neighbors and/or choose not to transmit their own features. This selective communication allows the model to focus on long-range dependencies and mitigate over-squashing by increasing the sensitivity of node features to information at distant nodes.
Motivated by these advances, we propose to combine the strengths of both paradigms through Cooperative Sheaf Neural Networks (CSNNs). CSNNs extend SNNs by incorporating the cooperative message-passing mechanism in the context of sheaves. To do so, we enrich undirected graphs into directed ones, define the corresponding notion of cellular sheaves, and introduce the first formulation of a sheaf Laplacian for directed graphs in machine learning. While this doubling of graph edges increases the number of restriction maps to be learned, potentially leading to overfitting, we address this by learning two restriction maps per node, in a similar fashion to Bamberger et al. (2024), instead of full sheaves. In our formulation, we assign conformal maps ${\mathbf{S}}_{i}$ and ${\mathbf{T}}_{i}$ to each node $i$ — corresponding to their role as a source and target in an edge, respectively — rather than learning separate restriction maps for each incident edge $i\unlhd e$. We define that a node $i$ does not broadcast to its neighbors when ${\mathbf{S}}_{i}=0$, and does not listen to its neighbors when ${\mathbf{T}}_{i}=0$. These maps are learned independently at each layer of the CSNN architecture.
Our theoretical and empirical results demonstrate the effectiveness of CSNNs in alleviating over-squashing, while leveraging the expressiveness and relational richness of both sheaf-based and cooperative frameworks. Our experiments show CSNN performs better than the prior sheaf-based models and cooperative GNNs in standard node classification benchmarks.
@@
    We provide a novel sheaf Laplacian over directed graphs, expanding the applications of sheaf theory in deep learning from undirected GNNs to directed GNNs.
@@
    For each layer $t$ in our CSNN model, each node is affected by information from nodes at distance up to $2t$-hop neighbors (Proposition 4.2), instead of up to $t$-hop neighbors in usual GNNs. Moreover, we can learn a sheaf in which the feature of a node $i$ at layer $t$ is highly sensitive to the initial feature of a node $j$, where $t$ is also the distance between $i$ and $j$ (Proposition 4.3).
@@
    We present a new SNN model that preserves its characteristics of avoiding oversmoothing and better perform on heterotrophic graphs, and additionally, mitigates overfitting and over-squashing, by using a similar approach proposed by Bamberger et al. (2024) but with an architecture closer to the NSD presented in Bodnar et al. (2022).
@@ 2 Background
We begin by introducing the core concepts of cellular sheaves over undirected graphs.
@@
Hereafter, we assume all vertex and edge stalks are isomorphic to ${\mathbb{R}}^{d}$. If all restriction maps are equal to the identity map, we say the cellular sheaf is constant. Furthermore, if $d\!=\!1$, the sheaf is said trivial.
Hansen and Ghrist (2021) interpret the node stalks ${\mathcal{F}}(i)$ as the space of private opinions of a person $i$. If $e$ is an edges between $i$ and $j$, then ${\mathcal{F}}(e)$ represents the space of public opinions discussed between these two nodes. We will leverage this allegory to provide intuitive accounts of technical definitions.
Given the importance of the Laplacian operator for graph representation learning, it is instrumental to define the Laplacian for undirected cellular sheaves – a key concept in the design of SNNs. Towards this end, the definition below introduces the domain and codomain of the coboundary operator.
@@
$C^{0}(G,{\mathcal{F}})=\bigoplus_{i\in V}{\mathcal{F}}(i)\mbox{ and }C^{1}(G,{\mathcal{F}})=\bigoplus_{e\in E}{\mathcal{F}}(e).$
@@
The Laplacian $L_{\mathcal{F}}$ can also be seen as a positive semidefinite matrix with diagonal blocks $L_{ii}=\sum_{i\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{i\unlhd e}$ and non-diagonal blocks $L_{ij}=L_{ij}^{\top}=-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{i\unlhd e}$.
@@
Note that $\ker L_{{\mathcal{F}}}=\{\mathbf{X}\in C^{0}(G,{\mathcal{F}})\>|\>\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\,\forall e=\{i,j\}\in E(G)\}$. Using the opinions analogy, this can be understood as the space of public agreement between all pairs of neighboring nodes $i$ and $j$. Note that $i$ and $j$ can have distinct opinions about the same topic on their respective private opinion spaces $\mathcal{F}(i)$ and $\mathcal{F}(j)$; however, when they publicly discuss about this topic, they may prefer to not manifest their true opinion. This apparent consensus lies in $\operatorname{Ker}L_{{\mathcal{F}}}$.
Vector bundles. When restriction maps are orthogonal, we call the sheaf a vector bundle. In this case, $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)$, for each vertex $i\in V$. Bodnar et al. (2022) found vector bundles lead to computationally more efficient NSD models and superior empirical performance.
Flat vector bundles. To further enhance the scalability of sheaf-based models, Bamberger et al. (2024) considers the special case of flat vector bundles in which to every node $i$ is assigned an orthogonal map ${\mathbf{O}}_{i}$, which replaces $\mathcal{F}_{i\unlhd e}$ for any adjacent edge $e$. This is a clear computational advantage since we just need to learn the orthogonal maps ${\mathbf{O}}_{i}$ instead of all restriction maps. In this case, $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-{\mathbf{O}}_{i}^{\top}{\mathbf{O}}_{j}\mathbf{x}_{j}\right)$, for each vertex $i\in V$.
@@ 3 Cellular sheaves for directed graphs
In this section, we construct a sheaf Laplacian for directed graphs. To the best of our knowledge, we are the first to discuss this in the machine learning community. Moreover, we have found a single paper proposing a (symmetric) sheaf Laplacian in a similar context (Sumray et al., 2024). Before deepening the discussion on the sheaf Laplacian, we introduce the definition of sheaves over directed graphs.
@@
Figure 1: On the left we have a cellular sheaf shown for a single edge of an undirected graph with stalks isomorphic to $\mathbb{R}^{2}$. The restriction maps $\mathcal{F}_{i\unlhd e},\mathcal{F}_{j\unlhd e}$ move the vector features between these spaces. On the right we have the analogous situation for a sheaf on a single pair of directed edges. Then there are four, possibly distinct, restriction maps $\mathcal{F}_{i\unlhd ij},\mathcal{F}_{i\unlhd ji},\mathcal{F}_{j\unlhd ij},\mathcal{F}_{j\unlhd ji}$.
Again, a common dimension $d$ is fixed for all node and edge stalks. Under the opinion analogy, directed sheaves allow for the restriction maps $\mathcal{F}_{i\unlhd ij}$ and $\mathcal{F}_{i\unlhd ji}$ to be different: we can have both edges $ij$ and $ji$ meaning that both people know each other but $j$ may not listen to $i$ $(\mathcal{F}_{i\unlhd ij}=0)$ or be less influenced by $i$ than $i$ is influenced by $j$, for instance.
For directed graphs, it is common to define both in- and out-degree Laplacians (Agaev and Chebotarev, 2005). Given a directed graph with possibly asymmetric adjacency matrix $A$, the out-degree Laplacian is $L^{\text{out}}:=D^{\text{out}}-A$ and the in-degree Laplacian is $L^{\text{in}}:=D^{\text{in}}-A$, with $D^{\text{in}}$ and $D^{\text{out}}$ denoting the diagonal matrices containing in- and out-degree of nodes in $V(G)$. Definition 3.2 generalizes these notions to directed sheaves. Note that if $(G,{\mathcal{F}})$ is the trivial sheaf, then $L^{\text{out}}_{\mathcal{F}}=(L^{\text{out}})^{\top}$ and $L^{\text{in}}_{\mathcal{F}}=L^{\text{in}}$.
@@
We anticipate that working directly with sheaves will lead to scalability issues and overfitting since a map would be learned for each node-edge pair, as reported in Bodnar et al. (2022) for undirected sheaf models. Bamberger et al. (2024) alleviate this issue using flat bundles: we only have to learn one orthogonal map to each node since ${\mathbf{O}}_{i}=\mathcal{F}_{i\unlhd e}$ for every adjacent $e$. Since the graphs are directed, we need to distinguish if the adjacent edge is pointing to $i$ or leaving $i$, then for each node-edge pair we learn two maps. More precisely, for each node $i$, we have the same source conformal map ${\mathbf{S}}_{i}=\mathcal{F}_{i\unlhd ij}$ and a target conformal map ${\mathbf{T}}_{i}=\mathcal{F}_{i\unlhd ji}$, for all neighbors $j$ of $i$.

### v2

@@ Abstract
Sheaf neural networks (SNNs) leverage cellular sheaves to induce flexible diffusion processes on graphs, generalizing the diffusion mechanism of classical graph neural networks. While SNNs have been shown to cope well with heterophilic tasks and alleviate oversmoothing, we show that there is further room for improving sheaf diffusion. More specifically, we argue that SNNs do not allow nodes to independently choose how they cooperate with their neighbors, i.e., whether they convey and/or gather information to/from their neighbors. To address this issue, we first introduce the notion of cellular sheaves over directed graphs and characterize their in- and out-degree Laplacians. We then leverage our construction to propose Cooperative Sheaf Neural Network (CSNN). Additionally, we formally characterize its receptive field and prove that it allows nodes to selectively attend (listen) to arbitrarily far nodes while ignoring all others in their path, which is key to alleviating oversquashing. Our results on synthetic data empirically substantiate our claims, showing that CSNN can handle long-range interactions without succumbing to oversquashing. We also show that CSNN performs strongly in heterophilic node classification and long-range graph classification benchmarks.
@@ 1 Introduction
Graph neural networks (GNNs) have become the standard models for an array of predictive tasks over networked data, with far-reaching applications in, e.g., physics simulation (Sanchez-Gonzalez et al., 2020), recommender systems (Ying et al., 2018), and molecular modeling (Duvenaud et al., 2015; Gilmer et al., 2017). Nonetheless, classical GNNs have well-known pitfalls. For instance, they typically struggle on heterophilic tasks (Zhu et al., 2020) — i.e., cases where connected nodes often belong to different classes or have dissimilar features. Furthermore, GNNs may also be susceptible to oversmoothing (Oono and Suzuki, 2020) and oversquashing (Alon and Yahav, 2021). Oversmoothing occurs when stacking multiple GNN layers yields increasingly similar node representations, whereas oversquashing refers to the loss of information when carrying information through increasingly long paths — due to the compression of exponentially growing information into fixed-size vectors.
A recent line of works (e.g., Hansen and Gebhart, 2020; Bodnar et al., 2022; Bamberger et al., 2025), which we henceforth refer to as Sheaf Neural Networks (SNNs), proposes modeling node interactions using cellular sheaves to achieve a principled solution to deal with oversmoothing and heterophilic tasks. A cellular sheaf $\mathcal{F}$} over an undirected graph associates (i) vector spaces ${\mathcal{F}}(i)$ and ${\mathcal{F}}(e)$, known as stalks, to each vertex $i$ and each edge $e$, and (ii) a linear map $\mathcal{F}_{i\unlhd e}$, known as a restriction map, to each incident vertex-edge pair $i\unlhd e$. These mathematical constructs induce a sheaf Laplacian which is governed by the restriction maps and generalizes the conventional Graph Laplacian.
In a parallel line of investigation, Finkelshtein et al. (2024) have recently shown that GNNs generally lack the flexibility to allow for nodes to individually select how they cooperate with their neighbors, i.e., choose whether they convey and/or gather information to/from their neighbors. This selective communication (also called cooperative behavior) is an especially desirable trait to tackle oversquashing, as it allows controlling the amount of information flowing between nodes. A natural question ensues: Can sheaf neural networks achieve cooperative behavior?
In this paper, we provide a negative answer to this question. More precisely, for SNNs to zero out all the incoming information at a node $i$, they must set $\mathcal{F}_{i\unlhd e}=0$ for all incident edges $e$, which also implies the information flowing from $i$ is suppressed (see Figure˜2). To circumvent this limitation, we introduce the notion of directed cellular sheaves (Definition˜3.2) and define their in- and out-degree Laplacians (Definition˜3.3). Leveraging these notions, we propose Cooperative Sheaf Neural Networks (CSNNs). Importantly, we show that cooperative behavior can be achieved using only a pair of restriction maps per node, which considerably increases computational efficiency compared to full sheaves — in which the amount of restriction maps increases linearly with the number of edges.
Our theoretical results show that CSNN allows for nodes to selectively listen to other arbitrarily distant nodes, which is a desirable trait to alleviate over-squashing. Our results on synthetic data specifically designed to induce over-squashing (Alon and Yahav, 2021) substantiate our claims, showcasing CSNNs’ superior potential to handle long-range dependencies. In addition, extensive real-world experiments on 11 node-classification benchmarks and 2 long-range graph-classification tasks demonstrate that CSNNs consistently outperform existing SNNs and cooperative GNNs.
@@
    We introduce the notions of in- and out-degree Laplacians for cellular sheaves over directed graphs, which can be used to model asymmetric relationships between nodes. We treat undirected edges as a pair of directed ones and leverage these constructions to propose CSNN — provably extending the flexibility of sheaf diffusion to accommodate cooperative behavior;
@@
    We provide a theoretical analysis of CSNN, showing that: ($a$) for each layer $t$ in CSNN, nodes may be affected by information from nodes at distance up to $2t$-hop neighbors (Proposition˜4.2), instead of up to $t$-hop neighbors in usual GNNs; and ($b$) there exist restriction maps which make the embedding of a node $i$ at layer $t$ highly sensitive to the initial feature of a node $j$, where $t$ is also the distance between $i$ and $j$ (Proposition˜4.3).
@@
    We carry an extensive experimental campaign to validate the effectiveness of CSNNs, encompassing both synthetic and real-world tasks. Our experiments on synthetic data show that CSNNs is remarkably capable of mitigating over-squashing and modeling long-range dependencies. Meanwhile, results on over 13 real-world tasks show CSNNs consistently surpass prior sheaf-based models and cooperative GNNs.
@@ 2 Background
For the sake of completeness, here we provide a summary of core concepts concerning cellular sheaves over undirected graphs. We also briefly discuss neural sheaf diffusion and cooperative GNNs.
In this work, we denote an undirected graph by a tuple $G=(V,E)$ where $V$ is a set of vertices (or nodes) and $E$ is a set of unordered pairs of (distinct) vertices, called edges, with $n=|V|$ and $m=|E|$. We denote the neighbors of a node $i$ in $G$ by $N(i)=\{j:\{i,j\}\in E\}$. Next, Definition˜2.1 introduces the notion of cellular sheaves over undirected graphs.
@@
Hereafter, we assume all vertex and edge stalks are isomorphic to ${\mathbb{R}}^{d}$. If all restriction maps are equal to the identity map, we say the cellular sheaf is constant. Moreover, if d=1d\!=\!1, the sheaf is said trivial.
Given the importance of the Laplacian operator for graph representation learning, it is instrumental to define the Laplacian for undirected cellular sheaves – a key concept in the design of SNNs. Towards this end, we introduce in Definition 2.2 the spaces of 0- and 1-cochains.
@@
$C^{0}(G,{\mathcal{F}})=\bigoplus_{i\in V}{\mathcal{F}}(i)\mbox{ and }C^{1}(G,{\mathcal{F}})=\bigoplus_{e\in E}{\mathcal{F}}(e).$ (1)
@@
The Laplacian $L_{\mathcal{F}}$ can also be seen as a positive semidefinite matrix with diagonal blocks $L_{ii}=\sum_{i\unlhd e}\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{i\unlhd e}$ and non-diagonal blocks $L_{ij}=L_{ij}^{\top}=-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}$.
To build intuition around Definition˜2.1, we may interpret the node stalks ${\mathcal{F}}(i)$ as the space of private opinions held by an individual $i$, following the perspective of Hansen and Ghrist (2021). For an edge $e$ connecting nodes $i$ and $j$, the stalk ${\mathcal{F}}(e)$ corresponds to the public opinions exchanged between them.
@@
That being said, note that $\ker L_{{\mathcal{F}}}=\{\mathbf{X}\in C^{0}(G,{\mathcal{F}})\>|\>\mathcal{F}_{i\unlhd e}\mathbf{x}_{i}=\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\,\forall e=\{i,j\}\in E(G)\}$. This can be understood as the space of public agreement between all pairs of neighboring nodes $i$ and $j$. Note that $i$ and $j$ can have distinct opinions about the same topic on their respective private opinion spaces $\mathcal{F}(i)$ and $\mathcal{F}(j)$; however, when they publicly discuss this topic, they may prefer to not manifest their true opinion. This apparent consensus lies in $\operatorname{Ker}L_{{\mathcal{F}}}$.
Vector bundles. When restriction maps are orthogonal, we call the sheaf a vector bundle. In this case, $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-\mathcal{F}_{i\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}\right)$, for any $i\in V$. Flat vector bundles are special cases of vector bundles in which we assign an orthogonal map ${\mathbf{O}}_{i}$ to each node $i$ and set $\mathcal{F}_{i\unlhd e}={\mathbf{O}}_{i}$ for all $e$ incident to $i$. This entails $L_{\mathcal{F}}(\mathbf{X})_{i}:=\sum_{i,j\unlhd e}\left(\mathbf{x}_{i}-{\mathbf{O}}_{i}^{\top}{\mathbf{O}}_{j}\mathbf{x}_{j}\right)$, for any $i\in V$. Note that flat vector bundles only comprise $n$ restriction maps as opposed to $2m$ maps in general cellular sheaves. Prior works (Bodnar et al., 2022; Bamberger et al., 2025) have leveraged these simpler constructions to propose computationally efficient sheaf-based neural networks.
Neural Sheaf Diffusion (NSD). Bodnar et al. (2022) introduce NSD building on Euler iterations of the heat equation induced by $\Delta_{{\mathcal{F}}}$, i.e., $\dot{\mathbf{X}}=-\Delta_{{\mathcal{F}}}\mathbf{X}$. First, we project the initial node features $\mathbf{X}$ into $h$ channels using an MLP $\mathbf{\eta}$, i.e., $\mathbf{X}_{0}=\mathbf{\eta}(\mathbf{X})\in{\mathbb{R}}^{nd\times h}$. Then, NSD recursively computes
$\mathbf{X}_{t+1}=(1+(\bm{1}_{n\times h}\otimes\varepsilon))\odot\mathbf{X}_{t}-\sigma(\Delta_{{\mathcal{F}}(t)}(\mathbf{I}\otimes\mathbf{W}_{1,t})\mathbf{X}_{t}\mathbf{W}_{2,t}),$ (3)
where $\bm{1}_{n\times h}$ is an $n\times h$ matrix of ones, $\Delta_{{\mathcal{F}}(t)}$ is the sheaf Laplacian at layer $t$, $\varepsilon\in[-1,1]^{d}$ is a (learned) vector scaling the features along each stalk dimension, $\sigma$ is an element-wise non-linearity, and ${\mathbf{W}}_{1,t}\in{\mathbb{R}}^{d\times d},{\mathbf{W}}_{2,t}\in{\mathbb{R}}^{h\times h}$ are weight matrices. Importantly, the restriction maps which govern $\Delta_{{\mathcal{F}}(t)}$ are learned in an end-to-end fashion, alongside ${\mathbf{W}}_{1,t}$ and ${\mathbf{W}}_{2,t}$.
Cooperative GNNs. Finkelshtein et al. (2024) recently proposed flexibilizing message-passing GNNs by treating nodes as players that can choose how they cooperate with their neighbors. More specifically, cooperative GNNs employ an auxiliary GNN, called the action network, that decides individually how each node partakes in the message passing of the base GNN, or environment network. The action network decides whether each node only propagates information (PROPAGATE), only gathers information from neighbors (LISTEN), does none (ISOLATE) or does both (STANDARD). Cooperative GNNs learn the action and environment GNNs simultaneously, using the straight-through Gumbel-Softmax estimator to propagate gradients through the discrete actions of the action network.
@@ 3 Cellular sheaves for directed graphs
We kick off this section addressing our initial research question: Can SNNs achieve cooperative behavior? Recall that communication between nodes in an SNN is governed by its sheaf Laplacian, which is induced by the restriction maps. Thus, in SNNs, picking a state among PROPAGATE, LISTEN, ISOLATE, and STANDARD node $i$ translates to choosing a suitable configuration for its respective restriction maps. The following result says that SNNs cannot fully alternate between these action.
Proposition 3.1.
Let $i\in V$. If $L_{{\mathcal{F}}}(\mathbf{X})_{i}$ does not depend on $\mathbf{x}_{j}$ for any $j\in V$ neighbor of $i$, then $L_{{\mathcal{F}}}(\mathbf{X})_{j}=0$ or $L_{\mathcal{F}}(\mathbf{X})_{j}=\sum_{j,i\unlhd e}\mathcal{F}_{j\unlhd e}^{\top}\mathcal{F}_{j\unlhd e}\mathbf{x}_{j}$.
Put plainly, Proposition˜3.1 states that the sheaf diffusion provides a framework where a node $i$ that does not LISTEN (since it does depend on $j$) must not PROPAGATE (since the update of $j$ does not depend on $i$), independently of the action that $j$ takes. In other words, PROPAGATE implies LISTEN, which means it collapses to ISOLATE (see Figure˜2).
We can circumvent this limitation by treating undirected edges as a pair of directed ones, creating an additional channel of communication between nodes. To accommodate directed edges (i.e., $E\subseteq V\times V$), we propose using cellular sheaves over directed graphs.
@@
Figure 1: On the left, a cellular sheaf shown for a single edge of an undirected graph with stalks isomorphic to $\mathbb{R}^{2}$. The restriction maps $\mathcal{F}_{i\unlhd e},\mathcal{F}_{j\unlhd e}$ move the vector features between these spaces. On the right, the analogous situation for a sheaf on a single pair of directed edges. Then there are four, possibly distinct, restriction maps $\mathcal{F}_{i\unlhd ij},\mathcal{F}_{i\unlhd ji},\mathcal{F}_{j\unlhd ij},\mathcal{F}_{j\unlhd ji}$.
For simplicity, again, we henceforth assume all node and edge stalks are $d$-dimensional.
We are now left with the task of defining sheaf Laplacians that can be used for information diffusion in directed graphs. For directed graphs, it is common to define both in- and out-degree Laplacians (Agaev and Chebotarev, 2005). Given a directed graph with possibly asymmetric adjacency matrix $A$, the out-degree Laplacian is $L^{\text{out}}:=D^{\text{out}}-A$ and the in-degree Laplacian is $L^{\text{in}}:=D^{\text{in}}-A$, with $D^{\text{in}}$ and $D^{\text{out}}$ denoting the diagonal matrices containing in- and out-degree of nodes in $V(G)$. Definition˜3.3 below generalizes these notions to sheaves over directed graphs.
@@
We note that if $(G,{\mathcal{F}})$ is the trivial sheaf, then $L^{\text{out}}_{\mathcal{F}}=(L^{\text{out}})^{\top}$ and $L^{\text{in}}_{\mathcal{F}}=L^{\text{in}}$. Flat vector bundles over directed graphs. We can also improve the parameter efficiency of cellular sheaves over directed graphs using flat vector bundles. Since the graphs are directed, we need to distinguish between edges with identical endpoints but with different orientations. Thus, for each node $i$, we assign a source conformal map ${\mathbf{S}}_{i}$ and a target conformal map ${\mathbf{T}}_{i}$, and set $\mathcal{F}_{i\unlhd ij}={\mathbf{S}}_{i}$ and $\mathcal{F}_{i\unlhd ji}={\mathbf{T}}_{i}$ for all neighbors $j$ of $i$.